solve by factoring and using the principle of zero products;
$x^2-9=0

Answer by ewatrrr(10682)


Hi
factoring
x^2-9=0
(x-3)(x+3) = 0 Note:SUM of the inner product(-3x) and the outer product(3x) = 0
Applying principle of zero products to solve for x
(x-3)=0
x = 3
(x+3) = 0
x = -3